OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation x^2/100+y^2/36= 1

OpenStudy (aum):

\[ \large \frac{x^2}{100} + \frac{y^2}{36} = 1 \\ \large \frac{x^2}{10^2} + \frac{y^2}{6^2} = 1 ~~~~ ----- (1)\\ \large \]Compare this to th general form\[ \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2} = 1 ~~~~ ----- (2) \\ \large \]where (h,k) is the center, a is the semi=major axis and b is the semi-minor axis. Can you find (h,k), a and b bu comparing (1) and (2)?

OpenStudy (anonymous):

a= 10 b=6 but how do you get h and k when you only have x and y without anything in front of it

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

x^2/100+y^2/36= 1, So x^2/10^2+y^2/6^2= 1, By this standard form of ellipse function, the center is (0,0); And a=10>b=6, Since a coordinates with x, so the foci are on the X axis, Since a^2=b^2+c^2, c=8, So foci are (-8,0) and (8,0) By definition of vertices, let y=0, so x=10 or -10 =|a|, so vertices are (10,0) and (-10,0)

OpenStudy (anonymous):

thank you so much @WilliamZhang !